Shift report 20000904 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Sep 04 2000 - 07:53:20 EDT

  • Next message: http account: "Shift report 20000904 08:00-16:00"

    ------------------------------------------------------------
    Date: Mon Sep  4 07:53:20 2000, Shift: 00:00-08:00
    Supervisor: MM, with: EJ
    Log book#5 pages 91 - 93
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 23:00 to 3:30 and again from 5:07 (not ended)
    2. Runs 2528, 2529, 2530 finished, MRS at90 degrees,d5=B 
    350A, FFS and BFS at 4 deg, all  magnets at A polarity.
    A total of 830k Trig 1 events. 
    Run 2531 is still running.
    
    3) T3 tripped in run 2530 probably in RHIC startup.
    No other new problems
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Sep 04 2000 - 07:53:58 EDT